A tunnel throught mountain is for a four-lane highway is to have semi-ellitical opening. The total width of the highway is to be 14.64m, and the height of the opening is 4.78 metres.
¡ Aproximately how wide the opening be?
¡¡ What will be the width of each lane if the barrier between the lanes is 0.6m?

Answer by greenestamps(13200)



I will guess that the problem is not presented correctly.

With the questions that are asked, the fact that the tunnel cross section is semi-elliptical is irrelevant....

Furthermore, with only the given information, the two lanes closest to the edge of the tunnel will have zero height on their edges, which is absurd....

Nevertheless... to answer the questions that are asked....

i. The width of the tunnel is given: 14.64m.

ii. With a total width of 14.64m, four lanes, and three barriers of 0.6m between lanes, the width available for the 4 lanes is 14.64-3(0.6) = 14.64-1.8 = 12.84m. That makes the width of each lane 12.84/4 = 3.21m.

Since the picture drawn by that response is nonsensical, I suggest you re-post the problem presenting the information clearly and correctly.